# Enhanced Feature Learning via Regularisation: Integrating Neural Networks and Kernel Methods

### v1

@@ A.4.3 Proof of Lemma 13
Proof [Lemma 13] Define $g_{\zeta}$ such that $g_{\zeta}(0)=0$ and $g_{\zeta}^{\prime}(x)=\min(|g^{\prime}(x)|,1/\zeta)\operatorname*{sign}(g^{\prime}(x))$. Note that $\|g_{\zeta}^{\prime}\|_{\infty}\leq\frac{1}{\zeta}$, thus $g_{\zeta}$ is $\frac{1}{\zeta}$-Lipschitz. Additionally, for any $a\in\mathbb{R}$,
@@
$\leq\sqrt{\int_{-\infty}^{+\infty}\mathds{1}_{|g^{\prime}(t)|\geq 1/\zeta}\,{\rm d}t\cdot\int_{-\infty}^{+\infty}\big{(}|g^{\prime}(t)|-1/\zeta\big{)}^{2}\,{\rm d}t}$
$\leq\sqrt{\int_{-\infty}^{+\infty}\mathds{1}_{|g^{\prime}(t)|\geq 1/\zeta}\,{\rm d}t}\quad{\rm since}\quad\int_{-\infty}^{+\infty}\big{(}g^{\prime}(t)\big{)}^{2}\,{\rm d}t\leq 1$
$\leq\zeta\quad{\rm since}\quad\int_{-\infty}^{+\infty}\mathds{1}_{|g^{\prime}(t)|\geq 1/\zeta}\frac{1}{\zeta^{2}}\,{\rm d}t\leq\int_{-\infty}^{+\infty}\mathds{1}_{|g^{\prime}(t)|\geq 1/\zeta}\big{(}g^{\prime}(t)\big{)}^{2}\,{\rm d}t\leq 1,$
@@
A.4.4 Proof of Lemma 17
Proof [Lemma 17] We begin with the bounded case. The bounds on the expectations are clearly valid. Then, since $1+\sqrt{\|X\|^{*}}$ is a bounded variable, it is necessarily subgaussian with a variance proxy bounded by $\frac{(1+\sqrt{R})^{2}}{2\log(2)}\leq(1+\sqrt{R})^{2}$ (Vershynin, 2018, Proposition 2.5.2 (iv)).
@@ A.5.1 Lemma 19 and its Proof
Proof [Lemma 19] By Bach (2024, Proposition 4.2), we have
@@ A.5.2 Lemma 20 and its Proof
Proof [Lemma 20] An equivalent definition of subgaussianity is that for all $\lambda\in\mathbb{R}$, if $6\sqrt{2e}\sigma|\lambda|\leq 1$, then $\mathbb{E}(e^{\lambda^{2}Z^{2}})\leq e^{(6\sqrt{2e})^{2}\sigma^{2}\lambda^{2}}$, see Vershynin (2018, Proposition 2.5.2).
@@ A.5.3 Lemma 21 and its Proof
Proof [Lemma 21] We use a specific version of McDiarmid’s inequality (Meir and Zhang, 2003, Theorem 3). First, we show that the conditions for applying the theorem are met. Let $\tilde{H}:=\{h:(x,y)\in\mathcal{X}\times\mathcal{Y}\to\ell(y,f(x))-\ell(y,\tilde{f}(x))\mid\Omega(f)\leq D,\Omega(\tilde{f})\leq D\}$. For any $\lambda>0$, we have
@@ B.1 Experiment 1: Optimisation procedure, Importance of Positive Homogeneous Kernel
In Experiment 6, BKRR and both versions of BKerNN had regularisation parameter fixed equal to $\max_{i\in[n]}(\|x_{i}\|_{2})/n$, where $n$ is the number of training samples (i.e. 400). Backtracking line search was used for BKerNN and the starting step-size was 500, while the number of iterations was 40. For ReLUNN, the batch size was 16, while the number of iterations was 2500 which corresponds to 100 epochs, and the step-size was set to 0.01.

### v2

@@ A.4.3 Proof of Lemma 13
Proof [Proof of Lemma 13] Define $g_{\zeta}$ such that $g_{\zeta}(0)=0$ and $g_{\zeta}^{\prime}(x)=\min(|g^{\prime}(x)|,1/\zeta)\operatorname*{sign}(g^{\prime}(x))$. Note that $\|g_{\zeta}^{\prime}\|_{\infty}\leq\frac{1}{\zeta}$, thus $g_{\zeta}$ is $\frac{1}{\zeta}$-Lipschitz. Additionally, for any $a\in\mathbb{R}$,
@@
${\color[rgb]{0,0,0}\leq\int_{-\infty}^{+\infty}\mathds{1}_{|g^{\prime}(t)|\geq 1/\zeta}|g^{\prime}(t)|\,{\rm d}t\leq\int_{-\infty}^{+\infty}\zeta|g^{\prime}(t)|^{2}\,{\rm d}t}$
${\color[rgb]{0,0,0}\leq\zeta\quad{\rm since}\quad\int_{-\infty}^{+\infty}\big{(}g^{\prime}(t)\big{)}^{2}\,{\rm d}t\leq 1},$
@@
A.4.4 Lemma 19 and its Proof
Lemma 19 (Gaussian Complexity of Finite Set of Lipschitz Functions)
Let $h_{1},\ldots h_{M}$ be 1-Lipschitz functions from $\mathbb{R}$ to $\mathbb{R}$ and let $\varepsilon$ be a random centred Gaussian vector with identity covariance matrix. Then
$\mathbb{E}_{\varepsilon}\bigg{(}\sup_{h\in\{h_{1},\ldots,h_{M}\},w\in\mathcal{S}^{d-1}}$ $\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h(w^{\top}x_{i})\bigg{)}$
$\leq\mathbb{E}_{\varepsilon}\left(\bigg{\|}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}x_{i}\bigg{\|}^{*}+\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\sqrt{2\log M}\right).$
This lemma is inspired by Bartlett and Mendelson (2002).
Proof [Proof of Lemma 19] We use Slepian’s lemma (Ledoux and Talagrand, 1991, Corollary 3.14). For $h\in\{h_{1},\ldots h_{M}\},w\in\mathcal{S}^{d-1}$, let
All in all $\mathbb{E}_{\varepsilon}((X_{h,w}-X_{\tilde{h},\tilde{w}})^{2})\leq\mathbb{E}_{\varepsilon,\tilde{\varepsilon}}((Y_{h,w}-Y_{\tilde{h},\tilde{w}})^{2})$ therefore we can apply Slepian’s lemma and obtain
$\leq\mathbb{E}_{\varepsilon,\tilde{\varepsilon}}\left(\sup_{h\in\{h_{1},\ldots,h_{M}\},w\in\mathcal{S}^{d-1}}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}w^{\top}x_{i}+\sum_{j=1}^{M}\tilde{\varepsilon}_{j}\mathds{1}_{h=h_{j}}\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\right).$
$\leq\mathbb{E}_{\varepsilon,\tilde{\varepsilon}}\left(\sup_{w\in\mathcal{S}^{d-1}}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}w^{\top}x_{i}+\sup_{h\in\{h_{1},\ldots,h_{M}\}}\sum_{j=1}^{M}\tilde{\varepsilon}_{j}\mathds{1}_{h=h_{j}}\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\right)$
$\leq\mathbb{E}_{\varepsilon,\tilde{\varepsilon}}\left(\sup_{w\in\mathcal{S}^{d-1}}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}w^{\top}x_{i}+\sup_{j\in[M]}\tilde{\varepsilon}_{j}\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\right)$
$\leq\mathbb{E}_{\varepsilon}\left(\bigg{\|}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}x_{i}\bigg{\|}^{*}\right)+\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\sqrt{2\log M}.$
A.4.5 Proof of Lemma 17
Proof [Proof of Lemma 17] We begin with the bounded case. The bounds on the expectations are clearly valid. Then, since $1+\sqrt{\|X\|^{*}}$ is a bounded variable, it is necessarily subgaussian with a variance proxy bounded by $\frac{(1+\sqrt{R})^{2}}{2\log(2)}\leq(1+\sqrt{R})^{2}$ (Vershynin, 2018, Proposition 2.5.2 (iv)).
@@ A.5.1 Lemma 20 and its Proof
Proof [Proof of Lemma 20] By Bach (2024, Proposition 4.2), we have
@@ A.5.2 Lemma 21 and its Proof
Proof [Proof of Lemma 21] An equivalent definition of subgaussianity is that for all $\lambda\in\mathbb{R}$, if $6\sqrt{2e}\sigma|\lambda|\leq 1$, then $\mathbb{E}(e^{\lambda^{2}Z^{2}})\leq e^{(6\sqrt{2e})^{2}\sigma^{2}\lambda^{2}}$, see Vershynin (2018, Proposition 2.5.2).
@@ A.5.3 Lemma 22 and its Proof
Proof [Proof of Lemma 22] We use a specific version of McDiarmid’s inequality (Meir and Zhang, 2003, Theorem 3). First, we show that the conditions for applying the theorem are met. Let $\tilde{H}:=\{h:(x,y)\in\mathcal{X}\times\mathcal{Y}\to\ell(y,f(x))-\ell(y,\tilde{f}(x))\mid\Omega(f)\leq D,\Omega(\tilde{f})\leq D\}$. For any $\lambda>0$, we have
@@ B.1 Experiment 1: Optimisation procedure, Importance of Positive Homogeneous Kernel
In Experiment 6, BKRR and both versions of BKerNN had regularisation parameter fixed equal to $\max_{i\in[n]}(\|x_{i}\|_{2})/n$, where $n$ is the number of training samples (i.e. 400 except for “tecator” where it was 192, “semeion” where it was 80 and “pah” where it was 64. The ratio of samples in the test set compared to the total dataset was nonetheless constant equal to 1/5, as in the other datasets). Backtracking line search was used for BKerNN and the starting step-size was 500, while the number of iterations was 40 (except for “semeion” where it was 4. For ReLUNN, the batch size was 16, while the number of iterations was 2500 which corresponds to 100 epochs (except for “semeion” where it was 250, and the step-size was set to 0.01.